# FirstReferences

### v2

@@ Acknowledgements.
We acknowledge Diego Fallas Padilla for careful reading of the manuscript and helpful suggestions. This work was supported in part by the NOTS cluster operated by Rice University’s Center for Research Computing (CRC). G.P. acknowledges support from the Welch Foundation Award (grant no. C-2154), the Office of Naval Research Young Investigator Program (grant no. N00014-22-1-2282), the NSF CAREER Award (grant no. PHY-2144910), the Army Research Office (W911NF22C0012), and the Office of Naval Research (grant no. N00014-23-1-2665). We acknowledge that this material is based on work supported by the U.S Department of Energy, Office of Science, Office of Nuclear Physics under the Early Career Award (grant no. DE-SC0023806). The isotopes used in this research were supplied by the US Department of Energy Isotope Program, managed by the Office of Isotope R&\&&D and Production. H.P. acknowledges support from the NSF (grant no. PHY-2207283) and the Welch Foundation (grant no. C-1669). Work at the Center for Theoretical Biological Physics was supported by the NSF (grant no. PHY-2019745). J.N.O. was also supported by the NSF (grant no. PHY-2210291). P.G.W. was also supported by the D. R. Bullard-Welch Chair at Rice University (grant no. C-0016). R.Z. is a cofounder and chief executive officer of TAMOS Inc. G.P. is a cofounder and chief scientist of TAMOS Inc. The other authors declare no competing interests.
@@ Appendix A System calibration
As described in the main text, we have full control over the system parameters and the properties of the reservoirs. Prior to realizing the two-mode LVCM dynamics with the twelve laser tones, we independently calibrate each term of the Hamiltonian and the motional cooling of each vibrational mode using their corresponding laser configurations, following the procedure detailed in Ref. [20], with additional steps for the drives associated with the second vibrational mode. The following table lists the values of the system parameters and motional cooling rates used to realize the LVCM dynamics in the main text:
Reported Trapped-ion parameters [$2\pi\times$ kHz]
@@
[…]
@@ Appendix B Numerical simulations
We use a QuTiP-based Python package [55] to numerically obtain the LVCM dynamics for comparing with the experimental data and theoretical investigations. Due to experimental imperfections, we include additional decoherence processes to our simulations of Eq. (2), as follows:
@@
where the jump operator $\sigma_{y}$ and its corresponding rate $\gamma_{z}$ account for spin dephasing induced by laser power fluctuations in the rotated spin basis ($z\leftrightarrow y$), while the jump operators $c_{im}=a_{i}a_{i}^{\dagger}+a_{i}^{\dagger}a_{i}$ and their corresponding rates $\gamma_{im}$ consider the motional dephasing of the radial tilt modes due to trap frequency fluctuations [56, 20]. From the comparison between the numerical calculations and experimental data, we obtain $\gamma_{z}/2\pi=7$ Hz and $\gamma_{im}/2\pi=20$ Hz for all cases.
@@ C.1 Weak electronic coupling
Degenerate case $(\omega_{1}=\omega_{2}\equiv\omega)$ As shown by the red and blue curves in Fig. S2A, the transfer rates of the two-mode system are lower to those of the single-mode system around $\Delta E=\omega$, and they become larger at higher resonances ($\Delta E=m\omega,m\in\mathbb{N},m>1$) for $\gamma_{1}+\gamma_{2}=\gamma$. This can be understood by comparing the FGR formula of the single-mode and two-mode cases. At a low temperature ($\bar{n}_{i}\sim 0.01$), the initial population in both cases dominantly occupies the ground level of the donor well. For $\Delta E=m\omega,m\in\mathbb{N}$, the transfer rates are proportional to the wavefunction overlaps between the donor and acceptor states, given by ${\rm FC}_{\text{2M}}\approx\sum_{k_{1}=0}^{m}\sum_{k_{2}=0}^{m}\delta_{k_{1}+k_{2},m}|\langle 0_{-}\ket{k_{1+}}|^{2}|\langle 0_{-}\ket{k_{2+}}|^{2}$ in the degenerate two-mode case and ${\rm FC}_{\text{1M}}\approx|\langle 0_{-}\ket{m_{+}}|^{2}$ in the single-mode case, where we suppose $p_{n_{i-}}\equiv p_{n_{-}}\approx p_{0_{-}}\approx 1$ and $|\langle n_{i-}\ket{n_{i+}}|\equiv|\langle n_{-}\ket{n_{+}}|\leq 1$ for simplicity. Intuitively, the vibrational modes of the two-dimensional donor well can “share” the energy difference, increasing the likelihood of transitions to higher excited states of the two-dimensional acceptor well (for $m>1$) despite the lessened individual couplings from the smaller wavefunction overlaps as compared to the single-mode case. For example, for a well separation of $g_{i}/\omega_{i}=1$ between the donor and acceptor energy surfaces, the Franck-Condon factors of the two systems decrease with increasing $\Delta E=m\omega$ as follows: ${\rm FC}_{\text{2M}}=\{0.271,0.271,0.180,0.124\}$ and ${\rm FC}_{\text{1M}}=\{0.368,0.184,0.061,0.015\}$ for $m=\{1,2,3,4\}$, which explains the comparative features of the red and blue data points in Fig. S2A.
However, for increasing width $\gamma$, the transfer dependence on $\Delta E$ can be diminished at the expense of the transfer rates at resonances. As shown in the green curve of Fig. S2A, when we increase the dissipation rate on the slow mode, the widths of the two-mode transfer resonances grow despite the lowered peak values, which increases the transition probabilities for off-resonant processes, thus making excitation transfer more robust to the energy offset in the donor-acceptor system.
Non-degenerate case $(\omega_{1}>\omega_{2})$ The additional resonances, provided by the slow mode, can also increase the transfer robustness to $\Delta E$ (see Fig. S2B). Extra broadening of the resonances from increased $\gamma_{\rm eff}$ can further enhance this robustness. Meanwhile, the transfer rates at resonances associated with the fast mode ($\Delta E=m\omega_{1},m\in\mathbb{N}$) are evidently reduced compared to the single-mode system, differently from the degenerate case. This decrease can be explained by the significantly lowered overlaps of the wavefunctions $\left(\sum_{\begin{subarray}{c}n_{1-},n_{1+}\\ n_{2-},n_{2+}\end{subarray}}{\rm FC}_{\begin{subarray}{c}n_{1-},n_{1+}\\ n_{2-},n_{2+}\end{subarray}}\ll\sum_{\begin{subarray}{c}n_{-},n_{+}\end{subarray}}{\rm FC}_{\begin{subarray}{c}n_{-},n_{+}\end{subarray}}\right)$ due to the mismatch of the resonances between those associated with the fast vibrational mode and those provided by the slow vibrational modes ($\Delta E=m\omega_{1}\neq m\omega_{2}$).
@@ C.2 Weak vibronic coupling (VAET)
|$\ket{e_{\pm},\!n_{1},\!n_{2}}\!=\!\Biggl{(}\!$$\pm$ $\!\frac{\epsilon\pm\Delta E/2}{\sqrt{2\epsilon(\epsilon\!\pm\!\Delta E/2)}}\!\ket{\uparrow}\!+\!\frac{V}{\sqrt{2\epsilon(\epsilon\!\pm\!\Delta E/2)}}\!\ket{\downarrow}\!\Biggr{)}$ (9)
$\otimes$ |$\ket{n_{1}}\otimes\ket{n_{2}},$
@@
×$\times\left[n_{jI}\delta_{n_{jF},n_{jI}-1}\!+\!(n_{jI}+1)\delta_{n_{jF},n_{jI}+1}\right]{\rm L}(E_{IF},\gamma_{j}),$
where index $k\neq j$, specifically $(j,k)$ can only be either $(1,2)$ or $(2,1)$, and $E_{IF}\equiv E_{e_{+},n_{1I},n_{2I}}-E_{e_{-},n_{1F},n_{2F}}=2\epsilon+(n_{1I}-n_{1F})\omega_{1}+(n_{2I}-n_{2F})\omega_{2}$ is the energy difference between the initial |$\ket{e_{+},n_{1I},n_{2I}}$ and final |$\ket{e_{-},n_{1F},n_{2F}}$ eigenstates. The allowed transfers in Eq. (C.2) describe the processes in which the vibronic system exchanges single-phonon energy with the environment to enable the excitation transfer [14, 21].
@@
×$\;\;\;\times\;(n_{jI}+1)\delta_{n_{jF},n_{jI}+1}{\rm L}(E_{IF},\gamma_{j}).$
@@
For completeness, in the expression above, we include the negative energy input resonances ($02\epsilon<$ 0) even if they are suppressed in the case of low temperatures investigated in this work. There are two classes of two-phonon exchange processes here: single-mode exchange and double-mode exchange [30]. In the former type, two phonons from a specific vibrational mode participate in the exchange, while the latter class consists of processes in which one phonon from each mode contributes to the two-phonon-assisted processes.
@@ Appendix D Solution for VAET under NIBA
where, in the continuum limit, the couplings between the spin and each bosonic mode and the frequencies of the infinite bosonic bath are characterized by the spectral density $J\left(\omega\right)=\pi\sum_{k}{\lambda_{k}^{2}\delta(\omega-\omega_{k})}$. It has also been shown in Ref. [33] that if the spectral density takes the form:
@@
[…]
@@
[…]
@@
[…]
@@ Appendix E Interference effects in two-mode VAET
×$\quad\quad\quad\times\;{\rm L}(E_{\rm dual},\gamma_{1}+\gamma_{2}),$
@@
×$\quad\quad\quad\times\;{\rm L}(E_{\rm single}^{i},2\gamma_{i}),$
respectively. When $\omega_{1}=\omega_{2}\equiv\omega$, $g_{1}=g_{2}\equiv g$, and $\gamma_{1}=\gamma_{2}\equiv\gamma$, Eqs. (23) and (24) give $k_{T}^{(2)}(E_{\rm dual})=2\times k_{T}^{(2)}(E_{\rm single}^{i})$, explaining the enhancement at the second resonance of the degenerate two-mode VAET compared to its single-mode counterpart, as observed in Fig. 4A. Similarly, for $\omega_{1}>\omega_{2}$ and $g_{i}\ll\omega_{i}$, we get $k_{T}^{(2)}(E_{\rm dual})\!>\!k_{T}^{(2)}(E_{\rm single}^{i})$, as observed in Fig. 4C.
Figure S4: Transfer rates for different numbers of vibrational modes involved in CT. (A) Transfer rate spectra for the degenerate case ($\omega_{1}=\omega_{2}=\omega_{3}\equiv\omega$) with $(V,g_{i},\gamma_{i})=(0.2,1,0.02)\omega$ for $i=1,2,$ and $3$. (B) Transfer rate spectra for the non-degenerate case ($\omega_{1}\!>\!\omega_{2}\!>\!\omega_{3}$) with $(V,g_{1},\gamma_{1},\omega_{2},\omega_{3})=(0.125,1,0.025,0.625,0.375)\omega_{1}$, $(g_{2},\gamma_{2})=(1,0.020)\omega_{2}$, and $(g_{3},\gamma_{3})=(1,0.033)\omega_{3}$. In both cases, the solid curves represent the transfer rates calculated from Eq. (2) using the definition in Eq. (3) with the same additional decoherences used in the main text. The blue, red, and green curves are the numerical results of the CT systems with one vibrational mode ($i=1$ only), two vibrational modes ($i=1$ and $2$), and three vibrational modes ($i=1,2,$ and $3$), respectively.
@@ Appendix F Beyond two-mode LVCM
Here, we show how the conclusions of our work on the two-mode models can be straightforwardly extended to the three-mode models and presumably to multi-mode models with $i>3$ with Figs. S4 and S5. Similar to the degenerate two-mode CT process, the presence of the additional vibrational mode in the degenerate three-mode CT case increases the phase-space volume of the hybridized energy surfaces (now three-dimensional), which leads to a wider $\Delta E$ region with monotonically increasing transfer rates and sharper peaks at high-$\Delta E$ ($\geq 5\omega$) resonances corresponding to the release of the initially trapped population in the three-dimensional upper adiabatic states (see Fig. S4A). Meanwhile, as shown in Fig. S4B, when the vibrational energies are not all equal, the transfer rate spectrum features a smooth transfer profile similar to that of two-mode CT except for the slightly increased rates, which can be explained by the additional transfer channels provided by the third vibrational mode.
@@
Figure S5: Transfer rates for different numbers of vibrational modes involved in VAET. (A) Transfer rate spectra for the degenerate case ($\omega_{1}=\omega_{2}=\omega_{3}\equiv\omega$) with $(V,g_{i},\gamma_{i})=(0.2,0.2,0.04)\omega$ for $i=1,2,$ and $3$. (B) Transfer rate spectra for the non-degenerate case ($\omega_{1}>\omega_{3}>\omega_{2}$) with $(V,g_{1},\gamma_{1},\omega_{2},\omega_{3})=(0.2,0.2,0.04,0.4,0.6)\omega_{1}$, $(g_{2},\gamma_{2})=(0.2,0.04)\omega_{2}$, and $(g_{3},\gamma_{3})=(0.2,0.04)\omega_{3}$. In both cases, the solid curves represent the transfer rates calculated from Eq. (2) using the definition in Eq. (3) with the same additional decoherences used in the main text. The blue, red, and green curves are the numerical results of the VAET systems with one vibrational mode ($i=1$ only), two vibrational modes ($i=1$ and $2$), and three vibrational modes ($i=1,2,$ and $3$), respectively.

### v3

@@ Acknowledgements.
We acknowledge Diego Fallas Padilla for careful reading of the manuscript and helpful suggestions. This work was supported in part by the NOTS cluster operated by Rice University’s Center for Research Computing (CRC). G.P. acknowledges support from the Welch Foundation Award (grant no. C-2154), the Office of Naval Research Young Investigator Program (grant no. N00014-22-1-2282), the NSF CAREER Award (grant no. PHY-2144910), the Army Research Office (W911NF22C0012), and the Office of Naval Research (grant no. N00014-23-1-2665). We acknowledge that this material is based on work supported by the U.S Department of Energy, Office of Science, Office of Nuclear Physics under the Early Career Award (grant no. DE-SC0023806). The isotopes used in this research were supplied by the US Department of Energy Isotope Program, managed by the Office of Isotope R&\&D and Production. H.P. acknowledges support from the NSF (grant no. PHY-2513089) and the Welch Foundation (grant no. C-1669). Work at the Center for Theoretical Biological Physics was supported by the NSF (grant no. PHY-2019745). J.N.O. was also supported by the NSF (grant no. PHY-2210291). P.G.W. was also supported by the D. R. Bullard-Welch Chair at Rice University (grant no. C-0016). Author contributions: V.S., M.D.S., A.M., G.T., R.Z., and G.P. contributed to the experimental design, construction, data collection, and analysis of this experiment. V.S., M.Z., H.P., J.N.O., P.G.W., and G.P. contributed to the paper’s conceptualization and supporting theory and numerics. All authors contributed to the writing and revision of the manuscript. Competing interests: R.Z. is a cofounder and chief executive officer of TAMOS Inc. G.P. is a cofounder and chief scientist of TAMOS Inc. The other authors declare no competing interests. Data availability: The authors declare that the data supporting the findings of this study are available within the paper, its appendices, and upon request.
@@ Appendix A System calibration
As described in the main text, we have full control over the system parameters and the properties of the reservoirs. Prior to realizing the two-mode LVCM dynamics with the twelve laser tones, we independently calibrate each term of the Hamiltonian and the motional cooling of each vibrational mode using their corresponding laser configurations, following the procedure detailed in Ref. [20], with additional steps for the drives associated with the second vibrational mode. These additional steps include calibrations of $g_{2}$, $\omega_{2}$, $\gamma_{2}$, $\bar{n}_{2}$, and $\bar{n}_{0,2}$, while maintaining the pre-calibrated values of $\Delta E$, $V$, $g_{1}$, $\omega_{1}$, $\gamma_{1}$, $\bar{n}_{1}$, and $\bar{n}_{0,1}$ (relevant to single-mode systems). The following table lists the values of the system parameters and motional cooling rates used to realize the LVCM dynamics in the main text:
Reported Trapped-ion parameters $2\pi\times$ kHz]
@@
Figs. 3A-B  $2\text{--}28\;$  1\;1\;  5\;5\;  6\;6\;  0.18\;0.18\;  5\;5\;  5.5\;5.5\; $0.2$
Figs. 3C-D  $3\text{--}40\;$  1.1\;1.1\;  8\;8\;  8.23\;8.23\;  0.18\;0.18\;  3\;3\;  2.98\;2.98\; $0.08$
Figs. 4A-B  $2\text{--}40\;$  0.5\;0.5\;  8\;8\;  10.3\;10.3\;  0.12\;0.12\;  3\;3\;  1.82\;1.82\; $0.2$
Figs. 5A-B  $5\text{--}25\;$  2\;2\;  10\;10\;  2\;2\;  0.45\;0.45\;  10\;10\;  2.2\;2.2\; $0.18$
Figs. 5C-D  $5\text{--}25\;$  2\;2\;  12\;12\;  3.3\;3.3\;  0.34\;0.34\;  8\;8\;  2.1\;2.1\; $0.15$
@@ Appendix B Numerical simulations
We use a QuTiP-based Python package [58] to numerically obtain the LVCM dynamics for theoretical investigations and to compare with the experimental data. Due to experimental imperfections, we include additional decoherence processes to our simulations of Eq. (4), as follows:
@@
where the jump operator $\sigma_{y}$ and its corresponding rate $\gamma_{z}$ account for spin dephasing induced by laser power fluctuations in the rotated spin basis ($z\leftrightarrow y$), while the jump operators $c_{im}=a_{i}^{\dagger}a_{i}$ and their corresponding rates $\gamma_{im}$ consider the motional dephasing of the radial tilt modes due to trap frequency fluctuations [59, 20]. From the comparison between the numerical calculations and experimental data, we obtain $\gamma_{z}/2\pi=7$ Hz and $\gamma_{im}/2\pi=80$ Hz for all cases. These experimental imperfections reduce the resolution in $\Delta E$ and suppress the sharp (and narrow) peaks in the transfer rate spectra.
@@ C.1 Weak electronic coupling
[…]
However, for increasing width $\gamma_{\rm eff}$, the transfer dependence on $\Delta E$ can be diminished at the expense of the transfer rates at resonances. As shown in the green curve of Fig. S2A, when we increase the dissipation rate on the slow mode, the widths of the two-mode transfer resonances grow despite the lowered peak values, which increases the transition probabilities for off-resonant processes, thus making excitation transfer more robust to the energy offset in the donor-acceptor system.
Non-degenerate case $(\omega_{1}>\omega_{2})$ - The additional resonances, provided by the slow mode, can also increase the transfer robustness to $\Delta E$ (see Fig. S2B). Extra broadening of the resonances from increased $\gamma_{\rm eff}$ can further enhance this robustness. Meanwhile, the transfer rates at resonances associated with the fast mode ($\Delta E=m\omega_{1},m\in\mathbb{N}$) are evidently reduced compared to the single-mode system, differently from the degenerate case. This decrease can be explained by the significantly lowered overlaps of the wavefunctions $\left(\sum_{\begin{subarray}{c}n_{1-},n_{1+}\\ n_{2-},n_{2+}\end{subarray}}{\rm FC}_{\begin{subarray}{c}n_{1-},n_{1+}\\ n_{2-},n_{2+}\end{subarray}}\ll\sum_{\begin{subarray}{c}n_{-},n_{+}\end{subarray}}{\rm FC}_{\begin{subarray}{c}n_{-},n_{+}\end{subarray}}\right)$ due to the mismatch of the resonances between those associated with the fast vibrational mode and those provided by the slow vibrational mode ($\Delta E=m\omega_{1}\neq m\omega_{2}$).
@@ C.2 Weak vibronic coupling (VAET)
[…]
@@
$\times\left[n_{jI}\delta_{n_{jF},n_{jI}-1}\!+\!(n_{jI}+1)\delta_{n_{jF},n_{jI}+1}\right]{\rm L}(E_{IF},\gamma_{j}),$
[…]
@@
$\;\;\;\times\;(n_{jI}+1)\delta_{n_{jF},n_{jI}+1}{\rm L}(E_{IF},\gamma_{j}).$
@@
For completeness, in the expression above, we include the negative energy input resonances ($2\epsilon<0$) even if they are suppressed in the case of low temperatures investigated in this work. There are two classes of two-phonon exchange processes here: single-mode exchange and double-mode exchange [27]. In the former type, two phonons from a specific vibrational mode participate in the exchange, while the latter class consists of processes in which one phonon from each mode contributes to the two-phonon-assisted processes.
@@ Appendix D Solution for VAET under NIBA
where, in the continuum limit, the couplings between the spin and each bosonic mode and the frequencies of the infinite bosonic bath are characterized by the spectral density $J\left(\omega\right)=\pi\sum_{k}{\lambda_{k}^{2}\delta(\omega-\omega_{k})}$. It has been shown in Ref. [36] that if the spectral density takes the form:
@@
$\partial_{t}\langle\sigma_{z}\rangle$ $=$ $-2V^{2}\int_{0}^{t}{ds\langle\sigma_{z}\rangle\left(s\right)e^{-iE\left(t-s\right)}\langle e^{-iB\left(s\right)}e^{iB\left(t\right)}\rangle}$ (20)
@@
$Q_{1}\left(\tau\right)$ $=$ $\frac{1}{\pi}\int_{-\infty}^{\infty}{d\omega J(\omega)\sin{\left(\omega\tau\right)}/\omega^{2}},$
$Q_{2}\left(\tau\right)$ $=$ $\frac{1}{\pi}\int_{-\infty}^{\infty}{d\omega J(\omega)\left(\cos{\left(\omega\tau\right)}-1\right)\coth{\left(\beta\omega/2\right)}/\omega^{2}}.$
@@
$\partial_{t}\left\langle\sigma_{z}\right\rangle$ $=$ $\int_{0}^{t}dsf\left(t-s\right)\left\langle\sigma_{z}\right\rangle\left(s\right),$
@@ Appendix E Interference effects in two-mode VAET
$\quad\quad\quad\times\;{\rm L}(E_{\rm dual},\gamma_{1}+\gamma_{2}),$
@@
$\quad\quad\quad\times\;{\rm L}(E_{\rm single}^{i},2\gamma_{i}),$
respectively. When $\omega_{1}=\omega_{2}\equiv\omega$, $g_{1}=g_{2}\equiv g$, and $\gamma_{1}=\gamma_{2}\equiv\gamma$, Eqs. (25) and (26) give $k_{T}^{(2)}(E_{\rm dual})=2\times k_{T}^{(2)}(E_{\rm single}^{i})$, explaining the enhancement at the second resonance of the degenerate two-mode VAET compared to its single-mode counterpart, as observed in Fig. 5A. It also suggests that the dual-phonon process associated with the $\omega_{1}+\omega_{2}$ energy input accounts for approximately half of this enhancement. Similarly, for $\omega_{1}>\omega_{2}$ and $g_{i}\ll\omega_{i}$, we get $k_{T}^{(2)}(E_{\rm dual})\!>\!k_{T}^{(2)}(E_{\rm single}^{i})$, as observed in Fig. 5C.
Figure S4: Transfer rates for different numbers of vibrational modes involved in CT. (A) Transfer rate spectra for the degenerate case ($\omega_{1}=\omega_{2}=\omega_{3}\equiv\omega$) with $(V,g_{i},\gamma_{i})=(0.2,1,0.02)\omega$ for $i=1,2,$ and $3$. (B) Transfer rate spectra for the non-degenerate case ($\omega_{1}\!>\!\omega_{2}\!>\!\omega_{3}$) with $(V,g_{1},\gamma_{1},\omega_{2},\omega_{3})=(0.125,1,0.025,0.625,0.375)\omega_{1}$, $(g_{2},\gamma_{2})=(1,0.020)\omega_{2}$, and $(g_{3},\gamma_{3})=(1,0.033)\omega_{3}$. In both cases, the solid curves represent the transfer rates calculated from Eq. (4) using the definition in Eq. (5) with the same additional decoherences used in the main text. The blue, red, and green curves are the numerical results of the CT systems with one vibrational mode ($i=1$ only), two vibrational modes ($i=1$ and 22), and three vibrational modes ($i=1,2,$ and 33), respectively.
@@ Appendix F Beyond two-mode LVCM
Here, we show how the conclusions of our work on the two-mode models can be straightforwardly extended to the three-mode models and presumably to multi-mode models with $i>3$ with Figs. S4 and S5. Similar to the degenerate two-mode CT process, the presence of the additional vibrational mode in the degenerate three-mode CT case increases the number of state configurations on the hybridized energy surfaces (now three-dimensional), which leads to a wider $\Delta E$ region with monotonically increasing transfer rates and sharper peaks at high-$\Delta E$ ($\geq 5\omega$) resonances corresponding to the release of the initially trapped population in the three-dimensional upper adiabatic states (see Fig. S4A). Meanwhile, as shown in Fig. S4B, when the vibrational energies are not all equal, the transfer rate spectrum features a smooth transfer profile similar to that of two-mode CT except for the slightly increased rates, which can be explained by the additional transfer channels provided by the third vibrational mode.
@@
Figure S5: Transfer rates for different numbers of vibrational modes involved in VAET. (A) Transfer rate spectra for the degenerate case ($\omega_{1}=\omega_{2}=\omega_{3}\equiv\omega$) with $(V,g_{i},\gamma_{i})=(0.2,0.2,0.04)\omega$ for $i=1,2,$ and $3$. (B) Transfer rate spectra for the non-degenerate case ($\omega_{1}>\omega_{3}>\omega_{2}$) with $(V,g_{1},\gamma_{1},\omega_{2},\omega_{3})=(0.2,0.2,0.04,0.4,0.6)\omega_{1}$, $(g_{2},\gamma_{2})=(0.2,0.04)\omega_{2}$, and $(g_{3},\gamma_{3})=(0.2,0.04)\omega_{3}$. In both cases, the solid curves represent the transfer rates calculated from Eq. (4) using the definition in Eq. (5) with the same additional decoherences used in the main text. The blue, red, and green curves are the numerical results of the VAET systems with one vibrational mode ($i=1$ only), two vibrational modes ($i=1$ and 22), and three vibrational modes ($i=1,2,$ and 33), respectively.